# Can the 3 neutrino masses really be found using SN 1987A data?

### v7

@@ I Introduction
Thus, it is usually assumed that the spread in neutrino emission times for SN 1987A is large enough to disguise any tiny differences in neutrino travel times from which the neutrino mass could be found, Here (and elsewhere) we make no such assumption, and instead use the SN 1987A data themselves to see what they reveal about whether the masses can be found. As a result, the author rediscovered what Ramanath Cowsik had found in 1988 about the neutrino masses being possibly revealed in the SN 1987A data. (Ehrlich2011, ) As we shall see, the data suggests $m^{2}$ values for the three neutrino masses, all of which are much larger than thought possible, and one of which is a tachyon, i.e., $m^{2}<0.$ (Bilaniuk1962, ), While this paper is not the first time the author has made this claim, there is much new here, including (1) a much improved statistical analysis, (2) a demonstration that finding individual neutrino masses is possible, given the best current neutrino emission model of SN 1987A and (3) an invaluable clue from the KATRIN experiment’s reported failure to observe sterile neutrinos. (KATRIN2025, ). This absence dramatically strengthens the case that the three neutrino masses really can be found from the SN 1987A data, and it suggests a way KATRIN can redo its analysis that may yield a confirmation or rejection of the preposterous-sounding claim suggested in the title. The paper also suggests a way that oscillation experiments and especially the IceCube experiment might supply such a confirmation.
@@ IV Discussion
IV.1 Why not exclude the LSD data?
@@
Initially, the author mistakenly rejected the LSD data. This rejection was not because it provided evidence for a $m^{2}<0$ neutrino, but because the idea of monochromatic neutrinos from a supernova seemed preposterous. However since this issue and all the other reasons many people have for rejecting the LSD data can be answered, there is no valid reason not to keep LSD data, especially because with three active neutrinos the mandated number of straight lines about which clustering should occur is three. The burden of proof here should not be why not exclude this burst, but rather why not keep it – and there really is no compelling reason to ignore it.
IV.2 Why are the LSD data so bunched in time?
@@
If the LSD burst of 5 neutrinos is genuine, it may look strange in Fig. 1 that while the other data points lie all along their respective lines the five LSD neutrinos are all bunched together in time at t=484t=-484 min or -17 ksec. The three straight line result would be more believable perhaps if the five neutrinos yielded data points at five widely spaced points on the negatively sloped line. But recall that the $t<0$ data have a kilosecond time scale. If the five LSD data points were spread out along the LSD line and not bunched in time they never could have been distinguished from background because researchers were looking for a concentration of neutrino events in a seconds-long time interval, i.e., a burst. The fact that the 5 LSD neutrinos were recorded at -17 ksec does, of course, not mean they went backward in time, only that they travelled faster than light, assuming they started at $t=0$ along with all the other three neutrino bursts. Because of that bunching at the left end of the LSD line the uncertainty in the slope (and the value of $m^{2}$) is $18\%$ or $m^{2}=-192,000\pm 34,000eV^{2}/c^{4}$.
IV.3 Why do the LSD neutrinos have the same E?
@@
IV.4 Alternate explanations for the LSD neutrinos?
@@
IV.5 Best size of the horizontal error bars?
@@
IV.6 A simulation of SN 1987A neutrino emissions
In 2025 Bozza et al. have calculated the neutrino luminosity from SN 1987A using a new model they recently developed. (Bozza2025, ) They claim the analysis is in many ways more refined and accurate than previous ones with respect to the temporal structure of the neutrino luminosity. The functional form of their computed luminosity has 9 parameters that are determined by a least squares fit to the SN 1987A data, and their result is shown in Fig. 4. Note that the maximum neutrino emission occurs at about $T=100ms=0.1sec$ and the bulk of the emissions occur for $t<1$ second. Their best fit parameter values are all in accordance with current supernova simulations, and they find surprisingly that $95\%$ of the analyzed events (even those at $t>>1$ sec) are signal rather than background.
These Bozza et al. results buttress our case that the SN 1987A neutrinos were mostly emitted during a 0.5 to 1.0 sec time window, and that our use of $\pm 0.5$ sec for the optimal horizontal uncertainty in our least squares fit to three straight lines was justified. A 0.5 to 1.0 sec time window is apparently small enough not to badly degrade the fit of 37 neutrino data points to three straight lines. However, the use of as many as 9 free parameters in the Bozza spectral fit is a serious drawback, since with enough free parameters one can fit virtually any shape. In contrast, our fit to the $m^{2}>0$ neutrino spectrum used only two parameters, i.e., the two $m^{2}>0$ masses. Moreover, as long as the spread in neutrino emission times is small compared to the horizontal error bars used (as implied by Fig. 2), relativistic kinematics requires that neutrinos of a given mass lie on straight lines having a specific slope, and there is no similar reqirement in the Bozza analysis.
Figure 4: Computed temporal electron antineutrino luminosity L in units of $10^{53}erg/s$ for SN 1987A neutrinos from Bozza et al. (Bozza2025, ). Note the time scale change at $t=1$ seconds. The fraction of the integrated luminosity for $t<1$ seconds is only a bit more than half because of that scale change. Figure is used under creative commons license https://creativecommons.org/licenses/by/4.0
@@ IV.8 The $\sum m_{\nu}$ constraint: a mass or flavor sum?
An apparent conflict exists between the three masses from the SN 1987A data fit and cosmological data. That conflict with the upper limit on the sum of the 3 neutrino masses $\Sigma m_{k}$ can be avoided, if one of the tachyon masses is negative gravitationally. (Schwartz2018, ) Thus, while tachyons are originally defined as having an imaginary kinematic mass, theorists often assume their gravitational mass is negative not imaginary, meaning they have a repulsive gravitational effect. The current upper limit on the total sum of neutrino masses from cosmology has tightened to $\sum m_{\nu}<0.07-0.12eV/c^{2}$ (Shao2025, ), but some analyses of the data even favor a negative value for the sum of the neutrino masses. (Elbers2025, ) There is an obvious conflict here between the current small limit on the sum and the three large masses found in the SN 1987A data, which could be avoided provided the mass sum were over flavor state masses not mass states.
@@ IV.9 Consistency with direct mass experiments?
KATRIN recently reported finding no evidence for sterile neutrinos in its data based on the observed shape of the tritium beta decay spectrum. (KATRIN2025, ). It conducted this search by fitting its data to a sum of two spectra: that for a zero mass neutrino plus the spectrum for a sterile neutrino of mass $m_{4}$, which was assumed to take on various values and various weights $\alpha=sin^{2}2\theta.$ At each $(m_{4},\alpha)$ grid point, the $\chi^{2}$ function is minimized with respect to three free parameters. By examining the excluded regions of the plane based on the high $\chi^{2}$ KATRIN concluded there was no evidence for sterile neutrinos at a high confidence level in their data over the entire grid. This negative result was was in conflict with the Neutrino-4 group that had previously found evidence for a $2.70eV/c^{2}$ sterile neutrino in their oscillation experiment. Moreover, the conflict of KATRIN with the Neutrino-4 experiment was not a small one, since the data of the GALLEX, SAGE, and BEST experiments confirms the parameters of neutrino oscillations $\Delta m^{2}_{14}=7.3eV^{2}$ (or $m_{4}=2.70eV/c^{2}$ and $\sin^{2}2\theta_{14}=0.36)$ and the confidence level of the Neutrino-4 sterile neutrinos combined with the data of the GALLEX, SAGE, and BEST experiments is now $5.8\sigma.$ (Serebrov2024, )
IV.11 “The dog that didn’t bark” (the key section!)
It is possible that KATRIN failed to see any sterile neutrinos because they simply do not exist, but here we consider a much more interesting possibility. Neutrino - 4 and the other three experiments that saw the $2.70eV/c^{2}$ neutrino were doing a short baseline disappearance oscillation experiment, which observed the disappearance probability $P(\bar{\nu_{e}}\rightarrow\bar{\nu_{e}}).$ This type of experiment would work just as well if the disappearing neutrinos were disappearing into sterile or active neutrinos. However, KATRIN’s search was specifically looking for sterile neutrinos and it would not have worked for active ones as it was conducted. (KATRIN was not being careless here, they just assumed like everyone else that neutrinos with large mass would have to be sterile) Of course, if the main claim of this paper is correct, and if the $2.70eV/c^{2}$ neutrino is an active neutrino (as well as a sterile one - see Fig. 3), then KATRIN’s failure to see it is understandable because their search method was not appropriate for active neutrinos. Specifically, KATRIN should not have included in the fitted combined differential spectrum any zero mass contribution. That’s because the zero mass effective mass is made up of the three specific masses we found for SN 1987A including one with $m^{2}<0.$ So, they should have fit their data using a combined differential spectrum $R_{C}$ for the three masses $m_{1}=2.70eV/c^{2},m_{2}=21.6eV/c^{2}$ and $m_{3}^{2}=-192,000eV^{2}/c^{4}$ with three adjustible weights, $\alpha_{j},$ that is:

### v8

# Finding the 3 neutrino masses using a one-phase $\nu$ emission model for SN 1987A

@@ I Introduction
Thus, it is usually assumed that the spread in neutrino emission times for SN 1987A is large enough to disguise any tiny differences in neutrino travel times from which the neutrino mass could be found, Here (and elsewhere) we make no such assumption, and instead use the SN 1987A data themselves to see what they reveal about whether the masses can be found. As a result, the author rediscovered what Ramanath Cowsik had found in 1988 about the neutrino masses being possibly revealed in the SN 1987A data. (Ehrlich2011, ) As we shall see, the data suggests $m^{2}$ values for the three neutrino masses, all of which are much larger than thought possible, and one of which is a tachyon, i.e., $m^{2}<0.$ (Bilaniuk1962, ), While this paper is not the first time the author has made this claim, there is much new here, including (1) a much improved statistical analysis, (2) a demonstration that finding individual neutrino masses is possible, (3) an extensive critique of the standard two-phase neutrino emission model of SN 1987A and (4) an invaluable clue from the KATRIN experiment’s reported failure to observe sterile neutrinos. (KATRIN2025, ). This absence dramatically strengthens the case that the three neutrino masses really can be found from the SN 1987A data, and it suggests a way KATRIN can redo its analysis that may yield a confirmation of the preposterous-sounding claim suggested in the title.
@@ IV Discussion
IV.1 Neutrino emission models: 1, 2 or 3 phases?
Standard neutrino emission models for supernovae usually stipulate two phases: an intense brief accretion phase lasting less than a second followed by a $10+$ seconds long cooling phase. Often the accretion phase is divided into two, with a first part being an initial breakout burst lasting under 50 ms, as the shock wave breaks through the surface of the protoneutron star, resulting in a massive flash of electron neutrinos. For example, Bozza et al. have calculated the neutrino luminosity from SN 1987A using a two phase model. (Bozza2025, ) The functional form of their two-phase computed luminosity has 9 free parameters that are determined by a least squares fit to the SN 1987A data, and their computed luminosity is shown in Fig. 4. Note that the maximum neutrino emission occurs at about $T=100ms=0.1sec,$ and the bulk of the neutrino emissions occur for $t<1$ seconds (the accretion phase). In contrast to the Bozza et al. two-phase model, our fit to the SN 1987A data just involves a single brief neutrino accretion phase lasting at most 0.5 to 1.0 seconds, and no cooling phase. An emission lasting at most 0.5 to 1.0 seconds was seen to be small enough so as to qualify as ”near-simultaneeous” neutrino emissions, for the purpose of giving a good fit to three straight lines on a plot of $1/E^{2}$ versus time – one line for each neutrino mass.
If two supernova neutrino emission models both give good fits to the SN 1987A data (the only test yet available), we may ask which model is more likely to be closest to reality, simplicity being an important criterion. A two phase emission model is less simple than a one phase model, and a three phase model would be even less simple. Actually, contrary to its description, the Bozza model in fact is a three phase model not two, as they admit in their paper: “the presence of a third stage is required in order to give a smooth behavior of the observable neutrino distributions in the whole time range.” (Bozza2025, ) Further favoring a one-phase model for SN 1987A over two or three is the fact that its cooling was peculiar and unexpected because it originated from a blue supergiant, rather than a red supergiant. This fact would lead to faster adiabatic cooling and a fainter initial object, which challenged traditional models. Giving still further support for a one phase model, the use of as many as 9 free parameters in the Bozza two-phase model is a serious drawback, since with that many free parameters one can find a good fit to virtually any shape emissions spectrum. In contrast, to the Bozza model, our one phase fit to the $t>0$ neutrino spectrum used only two parameters, namely the two $m^{2}>0$ neutrino masses found from the slopes in Fig. 1. Thus, as long as the spread in neutrino emission times is small, relativistic kinematics requires that neutrinos of a given mass lie on straight lines having a specific slope. Of course, there is no similar basic physics requirement justifying the two phases of accretion and cooling.
The only real justification for having those two separate phases is because that is what SN 1987A data seems to show, but only if the neutrino masses are small enough to ignore the differences in their energy-dependent travel times. However, this conclusion requires use of a circular argument, since their neutrino emission time spectrum only mirrors the observed arrival time spectrum if the neutrino masses are very small, and the Bozza model (unlike our one-phase model) just assumes this is the case. A final reason to be skeptical about the Bozza model is its choice of neutrino data to include in its analysis, particularly for the cooling region, i.e., $13<t<30$ seconds. Kamiokande II showed approximately 40 data points in this time interval having $N_{hits}>20$ $(E_{\nu}>7.2MeV)$ (Hirata1988, ). Of those 40 points spread over the 17 second interval Bozza selected only four data points for inclusion in their analysis, based on their neutrino scattering directions. (Bozza2025, ) Clearly, the goodness of the Bozza two (three) phase fit would depend greatly on their selection criteria on which of the 40 data points to include and exclude.
Despite the advantages discussed in this section of the author’s single phase neutrino emissions model over the Bozza two (or three) phase model for $t>0$ neutrinos, his inclusion of the $t<0$ LSD neutrino burst undoubtedly will result in some skepticism.
IV.2 Why not exclude the LSD data?
@@
Initially, the author did reject the LSD data. This rejection was not because it provided evidence for a $m^{2}<0$ neutrino, but because the idea of monochromatic neutrinos from a supernova seemed preposterous. However since this issue and most of the other reasons many people have for rejecting the LSD data can be dealt with, there is no valid reason not to keep LSD data, especially because given three active neutrinos, the mandated number of straight lines about which clustering of data points in Fig. 1 should occur is three. The burden of proof here should not be why not exclude this LSD burst, but rather why not keep it – and there really is no compelling reason to ignore it.
IV.3 Why are the LSD data so bunched in time?
@@
If the LSD burst of 5 neutrinos is genuine, it may look strange in Fig. 1 that while the other data points lie all along their respective lines the five LSD neutrinos are all bunched together in time at t=484t=-484 min or -17 ksec. The three straight line result would be more believable perhaps if the five neutrinos yielded data points at five widely spaced points on the negatively sloped line. But recall that the $t<0$ data have a kilosecond time scale. If the five LSD data points were spread out along the LSD line and not bunched in time they never could have been distinguished from background because researchers were looking for a concentration of neutrino events in a seconds-long time interval, i.e., a burst. The fact that the 5 LSD neutrinos were recorded at -17 ksec does, of course, not mean they “went backward in time,” only that they travelled faster than light, assuming they started at $t=0$ along with all the other three neutrino bursts.
IV.4 Why do the LSD neutrinos have the same E?
@@
IV.5 Alternate explanations for the LSD neutrinos?
@@
IV.6 Best size of the horizontal error bars?
@@
Figure 4: Computed temporal electron antineutrino luminosity L in units of $10^{53}erg/s$ for SN 1987A neutrinos from Bozza et al. (Bozza2025, ), who use a two phase model. Bozza’s accretion phase luminosity is in red, and their cooling phase is in blue. Note the time scale change at $t=1$ seconds. Figure is used under creative commons license https://creativecommons.org/licenses/by/4.0
@@ IV.8 The $\sum m_{\nu}$ constraint: a mass or flavor sum?
An apparent conflict exists between the three masses from the SN 1987A data fit and cosmological data. That conflict with the upper limit on the sum of the 3 neutrino masses $\Sigma m_{k}$ can be avoided, if one of the tachyon masses is negative gravitationally. (Schwartz2018, ) Thus, while tachyons are originally defined as having an imaginary kinematic mass, theorists often assume their gravitational mass is negative not imaginary, meaning they have a repulsive gravitational effect. The current upper limit on the total sum of neutrino masses from cosmology has tightened to $\sum m_{\nu}<0.07-0.12eV/c^{2}$ (Shao2025, ), but some analyses of the data even favor a negative value for the sum of the neutrino masses. (Elbers2025, ) There is an obvious conflict here between the current small limit on the sum and the three large masses found from the SN 1987A data, which could be avoided provided the mass sum were over flavor state masses not mass states.
@@ IV.9 Consistency with direct mass experiments?
In 2025 KATRIN reported finding no evidence for sterile neutrinos based on the observed shape of the tritium beta decay spectrum. (KATRIN2025, ). It conducted this search by fitting its data to a sum of two spectra: that for a zero mass neutrino plus the spectrum for a sterile neutrino of mass $m_{4}$, which was assumed to take on various values and various weights $\alpha=sin^{2}2\theta.$ At each $(m_{4},\alpha)$ grid point, the $\chi^{2}$ function is minimized with respect to three free parameters. By examining the excluded regions of the plane based on the high $\chi^{2}$ KATRIN concluded there was no evidence for sterile neutrinos at a high confidence level in their data over the entire grid. This negative result was was in conflict with the Neutrino-4 group that had previously found evidence for a $2.70eV/c^{2}$ sterile neutrino in their oscillation experiment. Moreover, the conflict of KATRIN with the Neutrino-4 experiment was not a small one, since the data of the GALLEX, SAGE, and BEST experiments confirms the parameters of neutrino oscillations $\Delta m^{2}_{14}=7.3eV^{2}$ (or $m_{4}=2.70eV/c^{2}$ and $\sin^{2}2\theta_{14}=0.36)$ and the confidence level of the Neutrino-4 sterile neutrinos combined with the data of the GALLEX, SAGE, and BEST experiments is now $5.8\sigma.$ (Serebrov2024, )
IV.11 Why KATRIN failed to see sterile neutrinos?
It is possible that KATRIN failed to detect any sterile neutrinos because they do not exist, but here we consider a much more interesting possibility. Neutrino - 4 and the other three experiments that saw the $2.70eV/c^{2}$ neutrino were doing a short baseline disappearance oscillation experiment, which observed the disappearance probability $P(\bar{\nu_{e}}\rightarrow\bar{\nu_{e}}).$ This type of experiment would work just as well if the disappearing neutrinos were disappearing into sterile or active neutrinos. However, KATRIN’s search was specifically looking for sterile neutrinos and it would not have worked for active ones as it was conducted. (KATRIN was not being careless here, they just assumed like everyone else that neutrinos with large mass would have to be sterile) Of course, if the main claim of this paper is correct, and if the $2.70eV/c^{2}$ neutrino is an active neutrino (as well as a sterile one - see Fig. 3), then KATRIN’s failure to see it is understandable because their search method was not appropriate for active neutrinos. Specifically, KATRIN should not have included in the fitted combined differential spectrum any zero mass contribution. That’s because the zero mass effective mass is made up of the three specific masses we found for SN 1987A including one with $m^{2}<0.$ So, they should have fit their data using a combined differential spectrum $R_{C}$ for the three masses $m_{1}=2.70eV/c^{2},m_{2}=21.6eV/c^{2}$ and $m_{3}^{2}=-192,000eV^{2}/c^{4}$ with three adjustible weights, $\alpha_{j},$ that is: